maxloaf1:

Identify the simulations that correctly use numbers to represent a situation in which 55% of store shoppers already have store credit cards.

maxloaf1:

Ultrilliam:

@Vocaloid

maxloaf1:

@Vocaloid

Ultrilliam:

*scratches head* I would help if I knew how to here... Voca's response time is usually much faster then this, like practically instantly x.x

Ultrilliam:

If it's ok, would you wait until she comes?

maxloaf1:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Ultrilliam If it's ok, would you wait until she comes? \(\color{#0cbb34}{\text{End of Quote}}\) Yes, ill wait

maxloaf1:

@Ultrilliam how much longerrr? :/

Ultrilliam:

Good question, she rarely takes this long unless she's AFK, and all our other good math people are offline... Desperate times call for desperate measures @ThisGirlPretty I call upon you for help lol

maxloaf1:

ANYONE HELP PLEASE!!!

Ultrilliam:

*notices she's also offline* wut... She's never offline...

maxloaf1:

I'm just going to take a wild guess then :/

Ultrilliam:

She'll be on soon then knowing her, she checks in every 10-15 minutes

Ultrilliam:

For good measure I'll tag her alt @thatcreepyboy

Ultrilliam:

Sorry we couldn't be of more help as quick as usual, it's just midnight for most of our users :(

maxloaf1:

It's alright @Ultrilliam, I really appreciate you trying to help me

Ultrilliam:

I apologize again

ThisGirlPretty:

Never fear tgp is here!

Tranquility:

What is 55% of 40? Hint: 55/100 * 40 = ??

Tranquility:

And do the same for 55% of 20, 60, and 80. That is 55/100 * 20 = ?? 55/100 * 60 = ?? 55/100 * 80 = ??